Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of eachbill does Anna have?
Answer by adamchapman(301)

let the number of $5 bills be called 'a' and the number of $10 bills be called 'b'

there are 12 bills in total, so:
a+b=12 ............(1)
the total value of all the bills is $100. So:
5a+10b-100 ............(2)
If we rearrange equation (1), we can get b=12-a, and replacin b in equation (2) with what we have just found:
5a+10(12-a)=100
5a+120-10a=100
-5a=-20
a=20/5
a=4
now replace the 'a' in equation (1) to get
4+b=12
b=8
Therefore Anna has 4 $5 bills and 8 $10 bills.
I hope this helps.
Please visit my website at www.geocities.com/quibowibbler